OpenStudy (dmezzullo):

Updated Batting Averages are computed using the formula: A = quantity of B plus K all over G A = batting average B = times at bat multiplied by player’s seasonal average K = hits in the new game G = times at bat plus times at bat in a new game Solve A = quantity of B plus K all over G for K. Please show your work. (6 points) Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the nearest whole number.

OpenStudy (dmezzullo):

Batter A B G K George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (dmezzullo):

a is 0.319 row b is 10.7 row g is 46 row and k have to figure out.... can someone help me?

OpenStudy (anonymous):

omg i just did that hold on let me see if i still have it

OpenStudy (dmezzullo):

k

OpenStudy (dmezzullo):

@jazy

OpenStudy (anonymous):

here we go ) Solve A = quantity of B plus K all over G for K. A = ( B + K ) / G A * G = B + K K = A * G - B 2) Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Let A = batting average Let B = times at bat multiplied by player’s seasonal average Let G = times at bat plus times at bat in a new game Let K = hits in the new game It is hard to make out the table, but I will presume the following: Player: George A = 0.319 B = 10.7 G = 46 Solving for K: K = A * G - B K = (0.319)(46) - 10.7 K = 14.674 - 10.7 K = 3.974 Rounded to the nearest whole number, K = 4 ----> George's hit in the a new game is 4. Player: Jose: A = 0.320 B = 11.4 G = 45 Solving for K: K = A * G - B K = (0.320)(45) - 11.4 K = 14.4 - 11.4 K = 3 ----> Jose's hits in a new game is 3. Player Dale: A = 0.298 B = 12.3 G = 52 Solving for K: K = A * G - B K = (0.298)(52) - 12.3 K = 15.496 - 12.3 K = 3.196 Rounded to the nearest whole number, K = 3 ----> Dale's hit in the a new game is 3. Player Michael A = 0.337 B = 15.2 G = 48 Solving for K: K = A * G - B K = (0.337)(48) - 15.2 K = 16.176 - 15.2 K = 0.976 Rounded to the nearest whole number, K = 1 ----> Michael's hit in the a new game is 1.

OpenStudy (dmezzullo):

put it in easier terms

OpenStudy (dmezzullo):

plz

OpenStudy (anonymous):

u cant u have to have all of that to get a hundred

OpenStudy (dmezzullo):

wat goes in tha k boxes?

OpenStudy (dmezzullo):

@mikala1

OpenStudy (anonymous):

all of that is there that is going to get you a 100 just copy and past it and delet the arrows

OpenStudy (dmezzullo):

tyvm